# Multimodal RGB-HSI Feature Fusion with Patient-Aware Incremental Heuristic Meta-Learning for Oral Lesion Classification

### v1

@@ 5 Conclusion
We presented a multimodal paradigm for classifying oral lesions into four categories. Our method uses deep morphological embeddings, haemoglobin-sensitive spectral indications, texture descriptors, and demographic data. Each modality captures lesion appearance and clinical context uniquely. RGB pictures were converted to 31-band hyperspectral images. These extracted spectral-style biomarkers are not available in RGB imaging. We created tiny texture and spectral-shape characteristics to capture local structure and wavelength-dependent variations. An incremental heuristic meta-learner (IHML) was introduced to fuse calibrated base models using uncertainty-aware meta-features and patient-level posterior smoothing, improving prediction stability. On an unseen patient split, IHML achieved a macro F1 of 66.23% and an AUC-ROC of 84.45%, outperforming all baseline models and confirming the benefit of multimodal feature integration. Future work will explore end-to-end fusion architectures, transformer-based attention, and real hyperspectral acquisition for improved clinical translation.

### v2

@@ 5 Conclusion
We presented a multimodal fusion framework for oral lesion classification that integrates deep visual embeddings, spectrally enriched descriptors, and clinical priors. Our method uses deep morphological embeddings, haemoglobin-sensitive spectral indications, texture descriptors, and demographic data. Each modality captures lesion appearance and clinical context uniquely. RGB pictures were converted to 31-band hyperspectral images. These extracted spectral-style biomarkers are not available in RGB imaging. We created tiny texture and spectral-shape characteristics to capture local structure and wavelength-dependent variations. We proposed integrating various features into IHML. IHML combines calibrated base learners, uncertainty-aware meta-features, and patient-level smoothing. This strategy improves stability and consistency in patient forecasts. Our strategy was tested on an unobserved patient split. IHML has a macro F1-score of 66.23% and an AUC-ROC of 84.45%. Our findings beat all baseline models. Each feature block yielded considerable gains in the ablation research. Our findings demonstrate that incorporating deep, spectral, and clinical information enhances oral lesion categorization. Future work will explore end-to-end multimodal attention architectures with explicit uncertainty modeling, as well as validation on physically acquired hyperspectral datasets.